OpenStudy (anonymous):

Please I need help with -11\13x<-2\3

OpenStudy (anonymous):

(x - 2)^3 = (x - 2)(x - 2)(x - 2) = (x - 2)(x^2 - 4x + 4) = x^3 - 4x^2 + 4x - 2x^2 + 8x - 8 = x^3 - 6x^2 + 12x - 8 this was an example for another answer but this is the procedure to do